OpenStudy (anonymous):

Which equation can be used to find eighty-three percent of seven hundred forty-nine? eighty-three over seven hundred forty-nine equals x over one hundred x over one hundred equals seven hundred forty-nine over eighty-three x over eighty-three equals one hundred over seven hundred forty-nine eighty-three over one hundred equals x over seven hundred forty-nine

OpenStudy (anonymous):

\[\frac{ 83 }{ 749 }=\frac{ x }{ 100 }\]-----------\[\frac{ x }{ 100 }=\frac{ 749 }{ 83 }\]-----------\[\frac{ x }{ 83 }=\frac{ 100 }{ 749 }\]-----------\[\frac{ 83 }{ 100 }=\frac{ x }{ 749 }\]----------- Do you know how to solve proportions?

OpenStudy (anonymous):

If so, I'll give you a hint. 83% of 749 is 621.67

OpenStudy (anonymous):

@aaronalbrecht66

OpenStudy (anonymous):

OK thank you.

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

If not, it was D.

OpenStudy (anonymous):

Yes i did